for a population with a standard deviation of 12 a score of x=44 corresponds to z= -.5 what is the population mean?
Answer by jim_thompson5910(35256)


z = (x - mu)/sigma
-0.5 = (44 - mu)/12
12*(-0.5) = 12*((44 - mu)/12)
-6 = 44 - mu
-6 -44 = 44 - mu - 44
-50 = -mu
-mu = -50
-mu/(-1) = -50/(-1)
mu = 50

The population mean (mu) is 50